OpenStudy (anonymous):

sin^2Ө+tan^2Ө+cos^2Ө= a) sec^2Ө b) sinӨ c) tan^2 Ө d) cos^3 Ө

OpenStudy (anonymous):

\[\sin ^{2}\theta+\cos ^{2}\theta=1\]

OpenStudy (anonymous):

\[Sin ^{2}(\Theta)+Tan ^{2}(\Theta)+Cos^{2}(\Theta)=Tan^{2}(\Theta)+1=Sec^{2}(\Theta)\]

OpenStudy (anonymous):

\[1+\tan ^{2}\theta=\sec ^{2}\]

OpenStudy (anonymous):

\[\sec ^{2}\theta \]

OpenStudy (anonymous):

a is the answerrrrrr

OpenStudy (anonymous):

Thanks everybody!